OpenStudy (anonymous):

3(sin 2x − cos 2x) write expression only in sine. I got sin=-(√3)/3 cos= (√3)/3 Does anyone know what to do next?

OpenStudy (anonymous):

Is it: \[\sin^2x\] Or \[\sin2x?\]

OpenStudy (anonymous):

sin2x

OpenStudy (anonymous):

And how do you get that values for sin and cos??

OpenStudy (anonymous):

using the Pythagorean theorem

OpenStudy (anonymous):

Can you show me how you used here?

OpenStudy (anonymous):

wait i just figured it out =-= i did it wrong lol

OpenStudy (anonymous):

\[3(\sin2x - (1 - 2\sin^2x)) \implies 3(\sin2x - 1 + 2\sin^2x) \implies 6\sin^2x +3\sin2x - 3\]

OpenStudy (anonymous):

I have used this: \[\huge \color{green}{\cos2x = 1 - 2\sin^2x}\]

OpenStudy (anonymous):

oh ok thanks :)